Evaluate, the radical expression if possible.
√4/5

Answer by dolly(163)

Given sqrt(4/5)

= sqrt(4)/sqrt(5)
= 2 / Sqrt(5)
Thus √4/5 = 2/√5
= 2/sqrt(5)